OpenStudy (anonymous):

Solve the equation (Completing the Square) x^2+7x+6=0

hero (hero):

To complete the square, do the following: 1. Put equation in the form ax^2 + bx = c x^2 + 7x = -6 2. Add (b/2)^2 to both sides: (7/2)^2 = 49/4 x^2 + 7x + 49/4 = -6 + 49/4 3. Simplify the left side into a perfect square: (x + 7/2)^2 = -6 + 49/4 4. Simplify the right side of the equation: (x+7/2)^2 = -24/4 + 49/4 (x+7/2)^2 = 25/4 Take the square root of both sides: x + 7/2 = + or - 5/2 Isolate x: x = -5/2 - 7/2 = -12/2 = -6 x = 5/2 - 7/2 = -2/2 = -1 x = (-6,-1)

OpenStudy (anonymous):

Thank you!(: